OpenStudy (anonymous):

find dy/dx of (X^y.Y^x=1)

OpenStudy (anonymous):

\[x^y y^x = 1\]

OpenStudy (anonymous):

yes how to solve it.. please tell...

OpenStudy (anonymous):

take log on both sides \[\log(x^y) + \log(y^x) = 0\] \[xlog(y) = -ylog(x)\]

OpenStudy (anonymous):

but we have to find dy/dx NOT Log...

OpenStudy (anonymous):

lol wait

OpenStudy (anonymous):

You know implicit differentiation ?

OpenStudy (anonymous):

No i dont know...

OpenStudy (anonymous):

Differentiate this last equation w.r.t x \[\frac{dx}{dx}\log(y) + \frac{x dy}{y dx} = -\frac{dy}{dx} - \frac{y}{x}\]

OpenStudy (anonymous):

Solve for dy/dx from this equation !

OpenStudy (anonymous):

@Ghayyas.mubashir

OpenStudy (anonymous):

i cant.. tell me some eazy method..

OpenStudy (anonymous):

Here y can't be totally expressed in terms of x... so implicit differentiation is the only way :)

OpenStudy (anonymous):

i Actually dont know implicit..

OpenStudy (anonymous):

@Meepi @amistre64

OpenStudy (anonymous):

A quick and handy way to do this sort of differentiation... (although, the method described above is the basis for it)... define a function f(x,y)=((x^y)(y^x))-1 if you know partial derivatives then \[dy/dx=-(\delta(f)/\delta(x))/(\delta(f)/\delta(y))\] where \[\delta\] represents partial derivative of function with respect to the variable.

OpenStudy (anonymous):

@Ghayyas.mubashir - the method described by @vikrantg4 is the easiest and most basic of all the methods with which the problem can be solved

OpenStudy (anonymous):

i dont understand his method

OpenStudy (anonymous):

You know chain rule and product rules ?

OpenStudy (anonymous):

@Ghayyas.mubashir - whose method ?

OpenStudy (anonymous):

no i dont know..

OpenStudy (anonymous):

he's referring to my method @xavier123

OpenStudy (amistre64):

Ghayyas, if you dont know implicit differentiation yet, then you are really not prepared for this question

OpenStudy (anonymous):

Here is an into to implicit differenctiation @Ghayyas.mubashir If you an equation expressed "implicitly" and not explicitly, i.e. not given in the y =... form but rather in a form like x^2 + y^2 = 1 - 2xy or something like that, then the way differentiate is by taking the derivative term by term with respect to x. If a function was given in y = .. form, how would you differentiate? You would obviously differentiate y with respect to x, so the derivative of a function like y = x^3 would be:\[y=x^3 \rightarrow \frac{ dy }{ dx }=3x^2\] Notice how the "dy/dx" notation really means that you are taking the chance in y with respect to x? You do the exact same with implicit differentiation. You look at the equation, something like the one I gave earlier, x^2 + y^2 = 1 - 2xy, and you treat y as a function of x. So let's differentiate this:\[x^2+y^2 = 1-2xy \rightarrow 2x+2y \frac{ dy }{ dx }=-2y-2x \frac{ dy }{ dx }\] See how each term containing "x" was differentiated normally but each term containing "y" was differentiate using chain rule. Since we are treating y as a function of x, the x are differentiated the way you would differentiate them in an equation like y = x^2, but y, is a function of x, so using chain rule, we have to do 2y and then multiply it by dy/dx to show that y is a function changing with respect to x. And then you just rearrange and factor the equation to solve for dy/dx. And once again, your dy/dx will also be expressed implicitly, i.e. it will have both x and y in it.

OpenStudy (anonymous):

@genius12 - nice explanation :)